# TritonDFT: Automating DFT with a Multi-Agent Framework

### 02 Mar 2026 v1

@@ 1 Introduction
Third, despite extensive benchmarks like graduate-level materials-domain knowledge (Zaki et al., 2024; Mirza et al., 2024), key capabilities in end-to-end DFT workflows, including numerical accuracy, Pareto-optimality, HPC parallelization, and cost efficiency, remain unevaluated. We present DFTBench to evaluate these capabilities. DFTBench comprises 73 materials spanning 10 distinct types to ensure diversity in physical properties and computational complexity, with expert-curated convergence tests totaling over 500 CPU-hours to obtain Pareto-optimal parameter configurations and ground-truth calculation results.
@@ 2 Background and Related Work
Figure 2: Energy Deviation and Computational Cost Variations with different DFT Parameters for Silicon (Space Group $Fd\bar{3}m$).
@@
TritonDFT (Ours) General QE Usage (>10>10 Types) Physics + HPC (Pareto-aware) Open Database + Memory + Human Interact. 10 (Metal, Insulator, Superconductor, Topological, …) Expert Curated Calculation
@@ 4 DFTBench: Benchmarking DFT Accuracy and Efficiency with LLM Agents
Benchmark Statistics. DFTBench is an expert-curated benchmark suite comprising 73 unique crystalline materials, exhibiting diversity in two aspects:
Physical Diversity: DFTBench covers 10 distinct material categories, ranging from fundamental electronic phases (Metals, Insulators, Semiconductors) to complex functional and quantum materials (Superconductors, Topological insulators, Ferroelectrics). The material set contains 47 chemical elements, and 23 crystallographic space groups, spanning multiple electronic phases and magnetic ground states. Such diversity in composition, symmetry, and electronic structure requires the model to understand different material properties and precisely configure $\boldsymbol{\theta}_{\mathrm{phy}}$.
@@ 5.2 Accuracy and Trade-off Analysis
DFT Parameter Evaluation. We evaluate the model performance in setting parameters for structural relaxation(VC-relax). It is the most fundamental step that determines the quality of the relaxed geometry and subsequent energy evaluations. Table 2 summarizes the results. Pass Rate denotes the fraction of cases where the generated configuration satisfies $\Delta E<20,10,1$ meV/atom, compared to the most accurate results obtained with the most stringent configuration. Normalized Cost is measured as the ratio of computational cost relative to the Pareto-optimal configuration obtained by convergence test. Values closer to 1 indicate stronger capability in identifying Pareto-optimal configurations. We also report the Advanced Requirement Satisfaction Rate, corresponding to parameters required by more complex materials such as spin polarization, Hubbard $U$ and van der Waals corrections. In total, 39 materials in DFTBench require one or more such advanced parameters.
@@
Model VC-relax (Lattice) SCF (E/atom) Band (Gap) DOS (Fermi)
@@
In terms of cost efficiency, GPT 4o performs best, while Gemini 2.5 Pro strikes a more balanced trade-off between pass rate and cost efficiency across different types. We also observe that for materials such as superconductors and insulators, cost efficiency tends to be lower. Although these systems are easier to achieve numerical convergence, models still tend to adopt more conservative parameter configurations, leading to unnecessary computational cost.
End-to-End Workflow Evaluation. We report the mean relative error of final results for different task workflows, using the maximum relative deviation of lattice parameters for VC-relax and total energy per atom for SCF. For band calculations and DOS, we report mean absolute error, measured by band gap and Fermi level respectively, since many ground-truth band gaps and Fermi levels are zero, making relative error ill-defined. All results are compared against the most stringent human-configured results. Across all four tasks, stronger models achieve consistently high accuracy, demonstrating the reliability of LLM-based agents for such fundamental tasks in end-to-end DFT workflows.
@@ 6 Conclusion
Future Direction. Future development will address the observed limitations in modeling complex quantum states, such as magnetic materials, where current models achieve pass rates below 6%, by integrating specialized physics-informed reasoning modules. We aim to expand the framework’s modularity beyond Quantum Espresso to support diverse DFT solvers, enhancing portability across the materials science ecosystem. Additionally, we plan to integrate TritonDFT with autonomous wet-lab platforms for automated theoretical simulation and validation, to enable a low-cost, high-throughput in closed-loop discovery.
@@ A.1 Information of Materials
DFTBench comprises 73 distinct materials spanning 10 different material categories. All material information is summarized in Table LABEL:tab:benchmark_data.
@@ 15 Fe metal I4/mmm 1
21 FeSe superconductor P4/nmm 4
22 MgB2 superconductor P6/mmm 3
[…]
24 NbSe2 superconductor P63/mmc 6
[…]
26 Sn superconductor I41/amd 2
[…]
28 Co magnetic P63/mmc 2
29 Cr2O3 magnetic R3c 10
@@ 35 NiO magnetic P4/mmm 8
[…]
43 BNhex semiconductor P63/mmc 4
@@ 54 Na3Bi topological P63/mmc 8
58 BaTiO3 ferroelectric P4mm 5
59 GeTe ferroelectric R3m 2
60 HfO2 ferroelectric Pca21 12
61 LiTaO3 ferroelectric R3c 10
@@
63 PbTiO3 ferroelectric P4mm 5
64 BN piezoelectric P63mc 4
65 CdS piezoelectric P63mc 4
66 CdSe piezoelectric P63mc 4
67 GaN piezoelectric P63mc 4
68 ZnO piezoelectric P63mc 4
@@ B.1 Information of Materials
DFTBench comprises 73 distinct materials spanning 10 different material categories. Some materials belong to multiple categories (e.g., semiconductor, optical). All material information is summarized in Table LABEL:tab:benchmark_data.
@@ B.2 Question Templates
Appendix C Experimental Details
All density functional theory (DFT) calculations were performed using Quantum ESPRESSO. Structural relaxations were carried out with a total energy convergence threshold of $1.0\times 10^{-9}$ Ry and a force convergence threshold of $1.0\times 10^{-8}$ Ry/Bohr. For the self-consistent field (SCF) cycles, the electronic convergence threshold was set to $1.0\times 10^{-10}$ Ry, and a mixing parameter $\beta=0.7$ was employed to ensure stable convergence. The plane-wave kinetic energy cutoff and Monkhorst–Pack $k$-point grids were determined from systematic convergence tests, requiring the total energy to be converged within 1 meV per atom. The converged cutoff energy and $k$-point mesh were subsequently used for all SCF calculations reported in this work. For non-self-consistent field (NSCF) calculations, denser $k$-point meshes were adopted to obtain accurate electronic properties such as band structures and density of states. The $k$-point density for NSCF calculations was increased on a case-by-case basis depending on the specific property being evaluated. Complete convergence data, including cutoff energies, $k$-point grids, and total energies, are provided in the accompanying CSV file to ensure reproducibility.

### 05 Mar 2026 v2

@@ 1 Introduction
Third, despite extensive benchmarks like graduate-level materials-domain knowledge (Zaki et al., 2024; Mirza et al., 2024), key capabilities in end-to-end DFT workflows, including numerical accuracy, Pareto-optimality, HPC parallelization, and cost efficiency, remain unevaluated. We present DFTBench to evaluate these capabilities. DFTBench comprises 68 materials spanning 10 distinct types to ensure diversity in physical properties and computational complexity, with expert-curated convergence tests totaling over 500 CPU-hours to obtain Pareto-optimal parameter configurations and ground-truth calculation results.
@@ 2 Background and Related Work
Figure 2: Energy Deviation and Computational Cost Variations with different DFT Parameters for Silicon (Space Group $Fd\bar{3}m$). Computational cost is normalized as $C/C_{0}$, where $C$ denotes the actual execution time and $C_{0}$ corresponds to the execution time with K=6 and ecut=40.
@@
TritonDFT (Ours) General QE Usage (>10>10 Types) Physics + HPC (Pareto-aware) Open Database + Memory + Human Interact. 10 (Metal, Insulator, van der Waals, Topological, …) Expert Curated Calculation
@@ 4 DFTBench: Benchmarking DFT Accuracy and Efficiency with LLM Agents
Benchmark Statistics. DFTBench is an expert-curated benchmark suite comprising 68 unique crystalline materials, exhibiting diversity in two aspects:
Physical Diversity: DFTBench covers 10 distinct material categories, ranging from fundamental electronic phases (Metals, Insulators, Semiconductors) to complex functional and quantum materials (van der Waals, Topological insulators, Ferroelectrics). The material set contains 47 chemical elements, and 23 crystallographic space groups, spanning multiple electronic phases and magnetic ground states. Such diversity in composition, symmetry, and electronic structure requires the model to understand different material properties and precisely configure $\boldsymbol{\theta}_{\mathrm{phy}}$.
@@ 5.2 Accuracy and Trade-off Analysis
DFT Parameter Evaluation. We evaluate the model performance in setting parameters for structural relaxation(VC-relax). It is the most fundamental step that determines the quality of the relaxed geometry and subsequent energy evaluations. Table 2 summarizes the results. Pass Rate denotes the fraction of cases where the generated configuration satisfies $\Delta E<20,10,1$ meV/atom, compared to the most accurate results obtained with the most stringent configuration obtained by human-performed convergence tests. Normalized Cost is measured as the ratio of computational cost relative to the Pareto-optimal configuration obtained by human-performed convergence tests. Values closer to 1 indicate stronger capability in identifying Pareto-optimal configurations. We also report the Advanced Requirement Satisfaction Rate, corresponding to parameters required by more complex materials such as spin polarization, Hubbard $U$ and van der Waals corrections. In total, 39 materials in DFTBench require one or more such advanced parameters.
@@
Model VC-relax (Lattice) SCF (E/atom) Bandgap (eV) EF (eV)
@@
In terms of cost efficiency, GPT 4o performs best, while Gemini 2.5 Pro strikes a more balanced trade-off between pass rate and cost efficiency across different types. We also observe that for materials such as van der Waals and insulators, cost efficiency tends to be lower. Although these systems are easier to achieve numerical convergence, models still tend to adopt more conservative parameter configurations, leading to unnecessary computational cost.
End-to-End Workflow Evaluation. Table 3 reports the mean relative error of final results for different task workflows, using the maximum relative deviation of lattice parameters for VC-relax and total energy per atom for SCF. For bandgap calculations and DOS, we report mean absolute error, measured by band gap and Fermi level respectively, since many ground-truth band gaps and Fermi levels are zero, making relative error ill-defined. All results are compared against the most stringent human-configured results. Across all four tasks, stronger models achieve consistently high accuracy, demonstrating the reliability of LLM-based agents for such fundamental tasks in end-to-end DFT workflows.
@@ 6 Conclusion
Future Direction. Future development will address the observed limitations in modeling complex quantum states, such as magnetic materials, where current models achieve pass rates below 6%, by integrating specialized physics-informed reasoning modules. We aim to expand the framework’s modularity beyond Quantum Espresso to support diverse DFT solvers, enhancing portability across the materials science ecosystem. Additionally, we plan to integrate TritonDFT with autonomous experimental platforms for automated theoretical simulation and validation, to enable a low-cost, high-throughput in closed-loop discovery.
@@ A.1 Information of Materials
DFTBench comprises 68 distinct materials spanning 10 different material categories. All material information is summarized in Table LABEL:tab:benchmark_data.
@@ 4 MgF2 Insulator P42/mnm 6
[…]
8 SiO2 Insulator P3121 9
9 BN (hex) Insulator P63/mmc 4
[…]
12 AlN Insulator, piezoelectric P63mc 4
13 BeO Insulator, piezoelectric P63mc 4
14 BaTiO3 Insulator, Ferroelectric P4mm 5
15 HfO2 Insulator, Ferroelectric Pca21 12
[…]
17 LiTaO3 Insulator, Ferroelectric, Piezoelectric R3c 10
18 PbTiO3 Insulator, Ferroelectric, Piezoelectric P4mm 5
19 Cr2O3 Insulator, Magnetic R3c 10
20 FeO Insulator, Magnetic P4/mmm 8
21 MnO Insulator, Magnetic P4/mmm 8
22 NiO Insulator, Magnetic P4/mmm 8
@@ 31 Fe Metal I4/mmm 1
37 MgB2 Metal P6/mmm 3
@@ 42 Co Metal, Magnetic P63/mmc 2
46 FeSe Metal, van der Waals P4/nmm 4
47 NbSe2 Metal, van der Waals P63/mmc 6
@@
61 GeTe Semiconductor, Ferroelectric R3m 2
62 CdS Semiconductor, Piezoelectric P63mc 4
63 CdSe Semiconductor, Piezoelectric P63mc 4
64 GaN Semiconductor, Piezoelectric P63mc 4
65 ZnO Semiconductor, Piezoelectric P63mc 4
@@
67 Na3Bi Topological P63/mmc 8
[…]
@@ B.1 Information of Materials
DFTBench comprises 68 distinct materials spanning 10 different material categories. Some materials belong to multiple categories (e.g., semiconductor, optical). All material information is summarized in Table LABEL:tab:benchmark_data.
@@ B.2 Question Templates
Appendix C Computational details for human-performed DFT calculations
All density functional theory (DFT) calculations were performed using Quantum ESPRESSO. Structural relaxations were carried out with a total energy convergence threshold of $1.0\times 10^{-9}$ Ry and a force convergence threshold of $1.0\times 10^{-8}$ Ry/Bohr. For the self-consistent field (SCF) cycles, the electronic convergence threshold was set to $1.0\times 10^{-10}$ Ry, and a mixing parameter $\beta=0.7$ was employed to ensure stable convergence. The plane-wave kinetic energy cutoff and Monkhorst–Pack $k$-point grids were determined from systematic convergence tests, requiring the total energy to be converged within 1 meV per atom. The converged cutoff energy and $k$-point mesh were subsequently used for all subsequent SCF calculations. For non-self-consistent field (NSCF) calculations, denser $k$-point meshes were adopted to obtain accurate electronic properties such as band structures and density of states. The $k$-point density for NSCF calculations was increased on a case-by-case basis depending on the specific property being evaluated. Complete convergence data, including cutoff energies, $k$-point grids, and total energies, are provided in the accompanying CSV file to ensure reproducibility.